# Funny-Valen-Tine: Planning Solution Distribution Enhances Machine Abstract Reasoning Ability

### v2

@@ Abstract
The importance of visual abstruct reasoning problems in the field of image processing is beyond words. Both Bongard-Logo problems and Raven’s Progressive Matrices (RPM) belong to the domain of visual abstract reasoning tasks, with Bongard-Logo categorized as image clustering reasoning and RPM involving image progression pattern reasoning. This paper introduces a novel baseline model, Valen, which falls under the umbrella of probabilistic highlighting models. Valen demonstrates remarkable performance in solving both RPM and Bongard-Logo problems, offering a versatile solution for these reasoning tasks. Our investigation extends beyond the application of Valen, delving into the underlying mechanisms of probability-highlighting solvers when tackling visual abstract reasoning problems. We revisit how these solvers handle instances in tasks like RPM and Bongard-Logo, realizing that they approximate the solution to each reasoning problem instance as a distribution delineated by primary and auxiliary samples. This led us to propose that the learning objective of probability-highlighting solvers is not the distribution of correct solutions but rather a distribution defined by both primary and auxiliary samples. To bridge the discrepancies, we introduced the Tine method, an adversarial learning-based approach designed to assist Valen in estimating a solution distribution closer to the correct one. However, Tine’s adversarial framework also introduces issues such as unstable training. Reflecting on Tine, we propose modeling the sample distribution of reasoning problems as a mixture of Gaussian distributions, leading to the development of the Funny method. Funny effectively models the sample distribution within reasoning problems as a mixture of Gaussian distributions, enabling Valen to more directly capture the true form of the correct solution distribution. Furthermore, we designed the SBR method to model the distribution of progressive patterns representation as a mixture of Gaussian distributions. Overall, this paper posits that the key to enhancing the ability of solvers to address visual abstract reasoning problems lies in explicitly planning the distribution of solutions to conform more closely to the distribution of correct solution.
Index Terms:
Abstract reasoning, Raven’s Progressive Matrices, Bongard-Logo, Adversarial Learning, Gaussian Mixture Model.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Deep neural networks have made remarkable achievements in various fields, including computer vision[1, 2, 3], natural language processing[4, 5, 6], generative model[7, 8, 9], visual question answering[10, 11], and visual abstract reasoning[14, 12, 13].
Visual abstract reasoning plays a pivotal role in the field of image processing. It is not only one of the advanced tasks in image processing, requiring systems to recognize objects in images and understand the relationships between them for logical inference, but also a key driver for technological innovation and application expansion in image processing. By addressing visual abstract reasoning problems, new algorithms and techniques such as deep learning and graph convolutional networks can be explored and applied, thereby propelling extensive applications and developments in fields like autonomous driving, medical diagnosis, and security surveillance. However, this field still faces challenges such as insufficient and imbalanced data, as well as limited model interpretability. Future research will focus more on enhancing the robustness, real-time performance, and interpretability of algorithms to further elevate the potential and value of visual abstract reasoning technology.
@@ I-A RAVEN and PGM
RPM[14], or Raven’s Progressive Matrices, is a classic visual reasoning intelligence test typically presented as a 3x3 matrix. It requires test-takers to observe and analyze visual features such as shape, size, direction, position, and their relationships within the problem statement, identify patterns or rules, and select an answer from given option pool that completes the matrix. Due to its non-verbal nature and value in assessing general intelligence, RPM is widely used in psychology, education, neuroscience, and other fields, serving as a tool for evaluating intelligence levels and studying human cognitive processes, intellectual development, and brain function. With the advancement of AI technology, researchers are exploring computer algorithms to solve RPM problems, albeit facing challenges. RAVEN[15] and PGM[16] are classic examples of RPM problems, with Figure 1 illustrating these two types of problems respectively.
@@ I-B Bongrad-Logo
Distinct from the image progressive pattern reasoning problems like RPM, the Bongard problem [12, 13] belongs to the category of image clustering reasoning problems. It typically involves multiple images divided into primary and auxiliary groups, with images in the primary group sharing an abstract concept governed by a specific rule and those in the auxiliary group deviating from these rules. Solving Bongard problems requires deep learning algorithms to accurately classify ungrouped images. As a specific instance within abstract reasoning, the Bongard-Logo problem [13] poses a significant challenge due to its inferential complexity. Each case comprises 14 images: six from the primary group, six from the auxiliary group, and two for classification. The grouping criterion relies on geometric shapes and their arrangement. Figure 2 illustrates a Bongard-Logo case.
@@ II Related work
II-A RPM solvers and Bongard-Logo solvers
There are numerous excellent solvers for RPM and Bongard-Logo problems, among which CoPINet[19], LEN+teacher[18], DCNet[20], NCD[22], SCL[23], SAVIR-T[24], PrAE[25], NVSA[27], ALANS[26], RS-TRAN[28] and CRAB[29] are notable RPM solvers, while SBPDL[30], PMoC[31] and Triple-CFN[32] are outstanding Bongard-Logo solvers. Since the method proposed in this paper is relatively independent and does not have a strong dependency on these existing works, the details about them will not be elaborated further.
@@ III Methodology
Both Bongard-Logo problems and Raven’s Progressive Matrices (RPM) fall under the umbrella of visual abstract reasoning tasks. Specifically, Bongard-Logo problems can be categorized as image clustering reasoning, while RPM challenges involve image progression pattern reasoning.
In this paper, we designed a general solver, Valen, for both RPM and Bongard-Logo problems. This paper proposes that to enhance the ability of solvers to address visual abstract reasoning problems, the distribution of solutions provided by the solver should be explicitly planned to better conform to the distribution of correct solutions. Therefore, we developed three accompanying methods: Funny, Tine, and SBR. These three methods all aim to explicitly plan the distribution of solutions generated by Valen.
Here we clarify a definition: in this paper, “primary samples” refer to those that align with the core patterns in reasoning problems, like correct RPM options or primary group samples in Bongard-Logo instances. Conversely, “auxiliary samples” are those that deviate from these patterns, such as incorrect RPM options or auxiliary group samples in Bongard-Logo problems.
@@ IV Valen: a new baseline model for RPM and Bongard-Logo problems
In this section, we introduce Valen (Visual Abstraction Learning Network), our baseline model designed to address both RPM, an image progressive pattern reasoning problem, and Bongard-Logo, an image clustering reasoning problem.
@@ IV-A Valen targeting RPM problem
The RPM problem challenges participants to complete a 3×3 matrix of images by choosing answers from a pool of options. A correct answer is one that, when placed in the matrix, maintains a consistent and specific image progression pattern that can be deduced from the incomplete matrix, referred to as the problem statement[36].
This paper requires a baseline to support our reasoning work. This baseline is expected to determine the plausibility of an RPM matrix based on the progression patterns exhibited within it and assign a probability to reflect this plausibility.
In RPM problems, there exists a unique design where an incomplete RPM matrix (missing one image) can still convey a complete progression pattern. This design is essential as RPM serves as an assessment tool. Hence, this paper asserts that the reasoning process of a successful RPM problem solver should be schematically represented as illustrated in the Figure 3.
Figure 3: The reasoning process of a envisioned RPM problem solver
As seen in the figure, our envisioned RPM solver consists of two modules: the Progressive Pattern Extractor and the Pattern Consistency Analyzer. Additionally, our envisioned RPM solver can perform the following three reasoning processes: 1. The process indicated by the blue line: When the RPM problem statement is completed by an option from the option pool to form a full $3\times 3$ progression matrix, the solver can enumerate all $N$ types of incomplete matrix of this $3\times 3$ matrix. 2. The process denoted by the green line: For each enumerated incomplete matrix, the Progressive Pattern Extractor extracts the progression pattern, resulting in a total of $N$ sets of progression patterns. 3. The process marked by the orange line: The solver utilizes its Pattern Consistency Analyzer to evaluate the consistency of these $N$ sets of progression patterns and outputs a score to quantify this consistency. This score also serves as the score for the corresponding option. The aforementioned three processes will also constitute the procedure of Valen.
Since this paper provides a clear definition of the reasoning process for a successful RPM solver, our designed Valen is derived by refining these three processes. Specifically, our Valen also consists of three processes: representation extraction of matrix and enumeration of incomplete matrix, extraction of progressive patterns from incomplete matrix, and consistency analysis of progressive patterns, which correspond to the aforementioned three processes respectively. We will introduce these three refined processes one by one.
IV-A1 Representation extraction of matrix and enumeration of incomplete matrix
In this paper, this process in Valen is designed as the structure shown in Figure 4.
@@ IV-A1 Representation extraction of matrix and enumeration of incomplete matrix
As shown in figure 4, when the problem statement of the RPM task is completed by a specific option to form a $3\times 3$ progression matrix, denoted as $\{x_{i}|i\in[1,9]\}$, Valen utilizes a basic Vision Transformer (ViT) as a feature extractor to obtain multi-viewpoint representations of $\{x_{i}|i\in[1,9]\}$. These representations are denoted as $\{r_{i}|i\in[1,9]\}_{v}$, where $v\in[1,V]$ and $V$ represents the number of viewpoints, $v$ represents the index of the viewpoint, and $i$ represents the index of the image in the matrix. The subsequent processes in Valen will handle these multi-viewpoint representation matrices in parallel and equally. The method of processing RPM images into multi-viewpoint representations has been applied in previous works [28, 24].
Subsequently, these multi-viewpoint representations $\{r_{i}|i\in[1,9]\}_{v}|_{v=1}^{V}$ will be enumerated by Valen to generate all possible incomplete representations matrix. The enumeration results are illustrated in the Figure 5.
@@
We denote these enumerated multi-viewpoint incomplete representation matrices as $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$, where $n$ represents the index of the incomplete representation matrix, and $i$ still represents the index of the representation within the incomplete matrix. Notably, the last item of the enumeration is the matrix with a incomplete option position, denoted as $\{r_{9i}|n\in[1,9],i\in[1,8]\}_{v}|_{v=1}^{V}$.
IV-A2 Extraction of progressive patterns from incomplete matrix
@@
In the figure 6, we can observe that we use an MLP to extract all row and column progressive information from the incomplete representation matrix $\{r_{9i}|i\in[1,8]\}_{v}$. Subsequently, we process all four row and column progressive information using a standard Transformer-Encoder to obtain $M$ representations of progressive patterns. $M$ is a hyperparameter that can be freely set, and in this paper, it is set to 2. The impact of setting $M$ to other values on Valen is not explored in this paper. This asymmetric seq-to-seq implementation (four-to-$M$) has been introduced in related techniques of ViT[17]. After processing all the incomplete representation matrices $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$ using the standard Transformer-Encoder, we obtain all multi-viewpoint progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$.
@@ IV-A3 Consistency analysis of progressive patterns
In this process, Valen will evaluate the consistency among these multi-viewpoint progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$. The task of this consistency analysis will be accomplished by the Transformer-Decoder. The process of evaluating the consistency of the progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}$ from a specific viewpoint $v$ can be conducted using a Transformer-Decoder, as illustrated in Figure 7.
@@
In the figure 7, we can observe that the progressive pattern representations $\{P_{9m}|m\in[1,M]\}_{v}$ extracted from the incomplete representation matrix $\{r_{9i}|i\in[1,8]\}_{v}$ are used as queries, while the other progressive pattern representations $\{P_{nm}|n\in[1,8],m\in[1,M]\}_{v}$, which are extracted from the incomplete representation matrices $\{r_{ni}|n\in[1,8],i\in[1,9],i\neq n\}_{v}$, serve as key-value pairs input into the Transformer-Decoder to compute the output vector set. The $M$ vectors within the output vector set are then individually mapped to a score by a new MLP. These mapped $M$ scores are summed to obtain the consistency score of the progressive patterns under viewpoint $v$.
Subsequently, We calculate the average of the progressive pattern consistency scores obtained from all viewpoints to determine the overall viewpoint score for the corresponding option. Finally, we employ cross-entropy as the loss function to constrain the scores of all options according to the logic of the RPM task, in order to optimize the learnable parameters involved in the aforementioned three processes.

### v3

@@ Abstract
The importance of visual abstract reasoning problems in the field of image processing cannot be overstated. Both Bongard-Logo problems and Raven’s Progressive Matrices (RPM) belong to the domain of visual abstract reasoning tasks, with Bongard-Logo categorized as image clustering reasoning and RPM involving image progression pattern reasoning. This paper introduces a novel baseline model, Valen, which falls under the umbrella of probability-highlighting models. Valen demonstrates remarkable performance in solving both RPM and Bongard-Logo problems, offering a versatile solution for these reasoning tasks. Our investigation extends beyond the application of Valen, delving into the underlying mechanisms of probability-highlighting solvers. In revisiting how these solvers handle RPM and Bongard-Logo tasks, we realize that they approximate the solution to each reasoning problem as a distribution in which primary samples are compliant while auxiliary samples are not. This prompts us to propose that the learning objective of probability-highlighting solvers is not the distribution of correct solutions but rather one jointly delineated by primary and auxiliary samples. To bridge the discrepancies, we introduced the Tine method, an adversarial learning-based approach that helps Valen estimate a distribution close to that of the correct solutions. However, adversarial training in Tine suffers from instability. Motivated by this limitation, we model the sample distribution of reasoning problems as a mixture of Gaussian distributions, enabling Valen to capture the correct solution distribution more efficiently. This non-adversarial methodology leads to the development of the Funny method. Building on a similar Gaussian-mixture paradigm, we further propose the SBR method to plan the distribution of progressive pattern representations. Overall, this paper contends that the key to enhancing solvers’ ability to address visual abstract reasoning problems lies in explicitly planning the distribution of predicted solutions to approach the correct solution distribution. Codes are available in: https://github.com/Yuanbeiming/Funny-Valen-Tine-Planning-Solution-Distribution-Enhances-Machine-Abstract-Reasoning-Ability
@@ I Introduction
Deep neural networks have made remarkable achievements in various fields, including computer vision [1, 2, 3, 4], natural language processing [6, 7, 5, 8], generative models [9, 10, 11], visual question answering tasks [12], and visual abstract reasoning [15, 13, 14].
Visual abstract reasoning plays a pivotal role in image processing. This task not only requires systems to detect objects and understand their relationships for logical inference, but also drives technological innovation and application expansion. By tackling such reasoning problems, we can stimulate the development of algorithms and technologies such as deep learning and graph convolutional networks, thereby advancing autonomous driving, medical diagnosis, security surveillance, and related fields. However, the field still faces insufficient and imbalanced data, limited model performance, and poor interpretability. Even large language models (LLMs), which have delivered breakthroughs in many domains, exhibit a significant performance drop on visual abstract reasoning tasks. Future research should therefore prioritize enhancing the abstract reasoning capabilities of machine intelligence and actively use such problems to evaluate the level of machine intelligence.
@@ I-A RAVEN and PGM
RPM[15], or Raven’s Progressive Matrices, is a classic visual reasoning intelligence test typically presented as a $\times3$ matrix. Test-takers must observe and analyze visual features such as shape, size, direction, position, and their relations, identify the underlying rules, and select the missing element from the given option pool that completes the matrix. Thanks to its non-verbal format and proven utility in assessing general intelligence, RPM is widely used in psychology, education, and neuroscience as a tool for evaluating intellectual ability and investigating human cognitive development, brain function, and related processes. With advances in AI, researchers are now exploring algorithmic solutions to the RPM problem, though significant challenges remain RAVEN[16] and PGM[17] are classic RPM benchmarks; Figure 1 illustrates sample instances. Intuitively, given the remarkable success of LLMs in many domains, one might expect them to handle such problems effortlessly. However, a series of studies [20, 18, 19] has shown that LLMs fall far short on RPM tasks, underscoring the unique value of these benchmarks and indicating that abstract reasoning remains a scarce and essential capability for machine intelligence.
@@ I-B Bongrad-Logo
Distinct from RPM tasks, the Bongard problem [13, 14] belongs to image clustering [38] reasoning problems. It typically involves multiple images divided into primary and auxiliary groups: images in the primary group share an abstract concept governed by specific rules, whereas images in the auxiliary group violate these rules. Solving Bongard problems demands that deep learning algorithms accurately classify the remaining ungrouped images. As a specific instance within abstract reasoning, the Bongard-Logo problem [14] poses a significant challenge due to its inferential complexity. Each case comprises 14 images: six from the primary group, six from the auxiliary group, and two ungrouped images to be classified. The grouping criterion relies on geometric shapes and their arrangement. Figure 2 illustrates one such case.
@@
Bongard-logo problems are categorized into four sets: The Free-form Shape Test Set (FF) aims to evaluate a model’s ability to generalize to more complex free-form shape concepts. The Basic Shape Test Set (BA) assesses the model’s generalization capability to new combinations of basic shape concepts. The Combinatorial Abstract Shape Test Set (CM) tests the model’s abstract understanding of novel attribute combinations, where all individual attributes have been seen in the training set, but the combinations are entirely new. The Novel Abstract Shape Test Set (NV) evaluates the model’s ability to discover new abstract concepts, particularly by withholding one attribute and all its combinations to examine the model’s extrapolation capability. Solvers are trained on the training set and tested on these four sets to evaluate their intelligence levels from multiple dimensions.
@@ II Related work
II-A RPM solver
Discriminative models for image reasoning, such as CoPINet [22], LEN+teacher [23], and DCNet [24], offer diverse solutions, focusing on learning differences and latent rules. Connectionist models: SCL [26], SAVIR-T [27], and neural-symbolic systems: PrAE [28], NVSA [29], ALANS [30] incorporate various techniques to improve reasoning accuracy and interpretability. By combining effective methods, RS-TRAN [32] have achieved impressive results on the RPM problem. Triple-CFN implicitly extracts information about concepts and reasoning units, and “indexes” this reasoning-unit information according to the concepts, achieving noteworthy reasoning accuracy. CRAB [33], based on a Bayesian modelling approach, has established a highly customised, controlled sandbox that aligns seamlessly with its proprietary methodology, embodied in the creation of a new RAVEN database. This bespoke environment, although it forfeits inherent core challenges of RAVEN such as solution diversity and uncertainty, has enabled CRAB to achieve remarkable breakthroughs. The community looks forward to assessing the broader implications of this approach for future research. DRNet [34] employs parallel CNN and ViT to extract local and spatial features, and uses a learnable fusion module to model abstract rules explicitly, offering a concise new framework for visual abstract reasoning. NCD [31] converts the unsupervised RPM task into contrastive learning via pseudo labels, and employs negative answer sampling plus decentralized features to adaptively cancel noise and inter-problem centroid shifts, enabling automatic rule representation extraction without any annotation and offering a scalable new paradigm for unsupervised abstract reasoning.
II-B Bongard solver
Researchers have tackled Bongard problems along three main avenues: symbolic models, convolutional neural networks, and synthetic-data strategies. Depweg et al. combined a formal language with symbolic vocabulary and Bayesian inference, yet their method struggled with complex instances and scaled poorly [13]. Kharagorgiev and Yun pre-trained CNNs on image datasets for feature extraction [35]. Nie et al. experimented with CNN-based meta-learning, but results remained sub-optimal [14]. These approaches exhibit distinct strengths and weaknesses, underscoring the need for further research. Yuan proposed PMoC [36], a probabilistic model that measures how likely auxiliary group samples fit the main group’s distribution—an approach that has shown strong performance. Notably, Triple-CFN [37] has successfully applied a unified framework to both Bongard-Logo and RPM problems.
@@ III Methodology
Both Bongard-Logo problems and Raven’s Progressive Matrices (RPM) fall under the umbrella of visual abstract reasoning tasks. Specifically, Bongard-Logo problems can be categorized as image clustering reasoning, while RPM challenges involve image progression pattern reasoning. We designed a general solver, Valen, for both RPM and Bongard-Logo problems. Furthermore, to enhance a solver’s ability to tackle visual abstract reasoning tasks, we propose that its solution distribution should be explicitly planned to conform more closely to that of correct solutions. Therefore, we developed three methods to perform this explicit planning: Funny, Tine and SBR.
Here we clarify the definitions: “primary samples” are those conforming to the core patterns of reasoning problems (e.g., correct RPM options or primary group samples in Bongard-Logo instances); “auxiliary samples” are those that do not (e.g., incorrect RPM options or auxiliary group samples in Bongard-Logo problems).
@@ IV Valen: a new baseline model for RPM and Bongard-Logo problems
In this section, we introduce Valen (Visual Abstraction Learning Network), our baseline model designed to address both RPM, an image progression pattern reasoning problem, and Bongard-Logo, an image clustering reasoning problem.
@@ IV-A Valen targeting RPM problem
The RPM problem challenges participants to complete a $\times3$ image matrix by selecting an answer from an option pool. The correct option is the one that preserves the progression patterns reasoned from the incomplete matrix (i.e., the problem statement) [41]. To this end, we introduce a baseline named Valen: it assesses the plausibility of the completed $\times3$ matrix based on its internal progressive patterns and outputs a corresponding probability.
A defining feature of RPM problems is that an incomplete matrix (missing one image) still retains sufficient information to reason about the complete progressive patterns. Based on this property, we argue that the reasoning process of an effective RPM problem solver should align with the illustration in Figure 3.
Figure 3: The reasoning process of an effective RPM problem solver
As seen in the figure, our envisioned RPM solver consists of two modules: the Progressive Pattern Extractor and the Pattern Consistency Analyzer. Additionally, our envisioned RPM solver can perform the following three reasoning processes: 1. The process indicated by the blue line: When an option from the pool completes the RPM problem statement into a full $3\times 3$ progression matrix, the solver can enumerate all $N$ possible incomplete versions of the completed $\times3$ matrix. 2. The process denoted by the green line: For each enumerated incomplete matrix, the Progressive Pattern Extractor extracts its progressive pattern, yielding $N$ sets of progressive patterns. 3. The process marked by the orange line: The solver utilizes its Pattern Consistency Analyzer to evaluate the consistency of these $N$ sets of progressive patterns and outputs a score to quantify this consistency. This score also serves as the score for the corresponding option.
Since this paper provides a clear definition of the reasoning process for a successful RPM solver, the Valen model is derived by refining these three steps. Specifically, Valen comprises three stages: (i) representation extraction of matrices and enumeration of incomplete matrices, (ii) extraction of progressive patterns from incomplete matrices, and (iii) consistency analysis of progressive patterns. We describe each stage individually.
IV-A1 Representation extraction of matrices and enumeration of incomplete matrices
This process in Valen follows the structure illustrated in Figure 4.
@@ IV-A1 Representation extraction of matrices and enumeration of incomplete matrices
As shown in this figure, when the problem statement of the RPM task is completed by a specific option to form a $3\times 3$ progression matrix, denoted as $\{x_{i}|i\in[1,9]\}$, Valen utilizes a basic Vision Transformer (ViT) as a feature extractor to obtain multi-viewpoint representations of $\{x_{i}|i\in[1,9]\}$. These representations are denoted as $\{r_{i}|i\in[1,9]\}_{v}$, where $v\in[1,V]$, $V$ is the total number of viewpoints, $v$ represents the index of the viewpoint, and $i$ represents the index of the image in the matrix. The subsequent processes in Valen handle these multi-viewpoint representation matrices in parallel and on an equal footing. The method of processing RPM images into multi-viewpoint representations has been applied in previous works [32, 27].
Subsequently, these multi-viewpoint representations $\{r_{i}|i\in[1,9]\}_{v}|_{v=1}^{V}$ will be enumerated by Valen to generate all possible incomplete representation matrices. The enumeration results are illustrated in Figure 5.
@@
We denote these enumerated multi-viewpoint incomplete representation matrices as $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$, where $n$ represents the index of the incomplete representation matrix, and $i$ still represents the index of the representation within the incomplete matrix. Notably, the last item of the enumeration is the matrix with an incomplete option position, denoted as $\{r_{9i}|i\in[1,8]\}_{v}|_{v=1}^{V}$.
IV-A2 Extraction of progressive patterns from incomplete matrices
@@
In the Figure 6, we can observe that we use an MLP to extract all row and column progressive information from the incomplete representation matrix $\{r_{9i}|i\in[1,8]\}_{v}$. Subsequently, we process all four row and column progressive information using a standard Transformer-Encoder to obtain $M$ representations of progressive patterns. $M$ is a hyperparameter that can be freely set, and in this paper, it is set to 2. The impact of setting $M$ to other values on Valen is not explored in this paper. This asymmetric seq-to-seq implementation (four-to-$M$) has been introduced in related techniques of ViT[21]. After processing all the incomplete representation matrices $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$ using the Transformer-Encoder, we obtain all multi-viewpoint progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$.
@@ IV-A3 Consistency analysis of progressive patterns
In this process, Valen evaluates the consistency among these multi-viewpoint progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$. As illustrated in Figure 7, Valen conducts this consistency evaluation among the representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}$ from each specific viewpoint $v$ using a Transformer-Decoder.
@@
In the Figure 7, the progressive pattern representations $\{P_{9m}|m\in[1,M]\}_{v}$ extracted from the incomplete representation matrix $\{r_{9i}|i\in[1,8]\}_{v}$ are used as queries, while the remaining progressive pattern representations $\{P_{nm}|n\in[1,8],m\in[1,M]\}_{v}$, which are extracted from the incomplete representation matrices $\{r_{ni}|n\in[1,8],i\in[1,9],i\neq n\}_{v}$, act as key-value pairs for the Transformer-Decoder to compute the output vector set. The $M$ output vectors are then individually mapped to scores by an MLP and summed to yield the consistency score of the progressive patterns under viewpoint $v$.
Subsequently, we compute the average of the progressive pattern consistency scores across all viewpoints to obtain the overall viewpoint score for the corresponding option. Finally, cross-entropy loss is employed to align the score distribution with the ground-truth option, thereby optimizing all learnable parameters involved in the aforementioned three processes.